# Prethermalization, shadowing breakdown, and the absence of Trotterization transition in quantum circuits

### v1

@@ Abstract
One of premier utilities of present day noisy quantum computers is simulation of many-body quantum systems. We study how long in time is such a discrete-time simulation representative of a continuous time Hamiltonian evolution, namely, a finite time-step introduces so-called Trotterization errors. We show that the truncated operator propagator (Ruelle-Pollicott resonances) is a powerful tool to that end, as well as to study prethermalization and discrete time crystals, including finding those phenomena at large gate duration, and can be used to calculate diffusion constant. We show that the effective energy is more stable than suggested by Trotter errors – a manifestation of prethermalization – while all other observables are not. Even such the most stable observable though deteriorates in the thermodynamic limit. Different than in classical systems with the strongest chaos, where the faithfulness time (the shadowing time) can be infinite, in quantum many-body chaotic systems it is finite. A corollary of our results is also that, opposite to previous claims, there is no Trotterization transition in non-integrable many-body quantum systems.
@@ I Introduction
One of the defining property of chaotic systems Ott is exponential complexity of their time evolution as quantified for instance by positive dynamical entropies. A manifestation of that is for instance the colloquial “butterfly effect” – two close initial conditions will separate exponentially fast in time. Considering that typical chaotic systems are not analytically solvable, and one therefore has to resort to the indignity of numerical simulations, a pertinent question is how representative are such noisy numerical trajectories of true trajectories? At first sight it would seem that the exponential separation of trajectories renders any noisy numerical simulation futile. However, things are not as grim.
First, we should understand better what is it that we really want from a noisy simulation. Comparing a noisy and a true trajectory both starting from the same specific initial condition is not what we should in most cases care about. This is especially true in a many-body system, where we do not have access to all degrees of freedom, and in quantum mechanics where the stress is on (ensemble) averages rather than on individual pure states. Calculating the overlap of two states, often called the fidelity (a.k.a. the Loschmidt echo), is not the most informative quantity in a many-body system (unless one is e.g. interested in the ground-state physics of a gapped system): small fidelity does not necessarily mean that few-body observables that we care about have gone awry. The properties of a specific initial state are really irrelevant; measurable quantities away from phase transitions should be robust to small variations. What one should care about is that the noisy evolution of a particular initial state is still representative, i.e., that its results are similar to a true ensemble average, or, at most, that the noisy evolution is similar to a true evolution from some nearby initial condition. Precisely that property is captured by the famous shadowing property of chaotic trajectories Ott that comes to our rescue when doing noisy simulations of chaotic systems.
@@
It states that while it is true that a trajectory $f^{t}_{\epsilon}(\mathbf{x}_{0})$ starting from $\mathbf{x}_{0}$ will in the presence of noise $\epsilon$ diverge exponentially fast from a true one $f^{t}_{0}(\mathbf{x}_{0})$ starting from $\mathbf{x}_{0}$, there exists a nearby initial condition $\mathbf{x}_{0}^{\prime}=\mathbf{x}_{0}+\bm{\delta}$ for which a true trajectory $f^{t}_{0}(\mathbf{x}_{0}^{\prime})$ closely follows the original noisy one, $f^{t}_{0}(\mathbf{x}_{0}^{\prime})\approx f^{t}_{\epsilon}(\mathbf{x}_{0})$ (Fig. 1). For hyperbolic systems the above intuitive picture can be made rigorous in the form of the shadowing lemma Anosov67 ; Bowen75 . Hyperbolic systems are roughly speaking systems where at each point there is an unstable (stretching) and a stable (folding) manifold. If a system is in addition uniformly hyperbolic, meaning that stretching exponents are uniformly larger than $1$ (i.e., stretching never becomes zero, or, stable and unstable directions never degenerate to coincide), then one can show that the time upto which the shadowing property holds is infinite. Those are in a way the strongest (nicest) chaotic systems, an example being the Arnold’s cat map. For classical systems that are not uniformly hyperbolic the shadowing time is finite Yorke97 .
We address the question of the shadowing time in quantum many body systems. This is an important theoretical question, as well as very relevant for noisy quantum computers and simulators, whose best-case present use troyer23 is to simulate daley22 other quantum systems ibm ; google ; dario ; shtanko25 ; feig25 , including the prethermalization regime banuls23 that we shall touch upon. If one wants to demonstrate the supremacy of quantum computation the number of qubits has to be large, and the simulation has to be faithful to true dynamics. Simulating many-body quantum Hamiltonian $H_{0}=\sum_{j}h_{j}$ with local interactions $h_{j}$ can be implemented by splitting the time evolution propagator
@@
Characterizing shadowing properties of a many-body quantum system is a formidable task. Here we tackle simpler question of direct experimental relevance that will nevertheless allow us to make conclusions also about the shadowing time. We will (i) focus on a specific type of noise, namely errors childs19 ; cubitt21 ; Heyl due to Trotterizing Hamiltonian evolution (2), and (ii) measure closeness by simple expectation values or autocorrelation functions of local observables (rather than some state norm).
@@
As we shall see, there is a good and a bad news. Trotterization errors are rather particular. A special observable, namely the energy $H_{0}$, will be very stable and resilient to errors, more than the simple Trotter error ${\cal O}(\tau^{2})$ would suggest, while other will be not. Despite its stability though, with increasing system size $L$ the stability will deteriorate. In the thermodynamic limit of $L\to\infty$ and time $t\to\infty$, no matter how small $\tau$, the energy $H_{0}$ will not be conserved implying that the noisy simulation is fundamentally different than the true one. As a consequence, perfect infinite-time shadowing, like the one holding for classical systems with the strongest chaos, is impossible under Trotterization errors in locally interacting quantum many-body systems. The relevant timescale of the above instability is nothing but the prethermalization Mori16 ; Abanin17 time $T$, for a review of prethermalization see Ref. ho23 , and is given by the largest eigenvalue of the momentum-resolved truncated operator propagator. As predicted theoretically, it is exponentially large in the gate frequency $1/\tau$, but the benefit of our approach will be to give an exact quantitative prediction for $T$, not just a bound. The corresponding eigenvector is on the other hand an almost conserved effective energy which is for small $\tau$ close to $H_{0}$.
@@
The truncated propagator therefore offers a way to obtain the effective Hamiltonian useful in studies of Floquet systems Eckardt17 , and is an explicit and concrete alternative to various divergent operator expansions like the Baker-Cambell-Hausdorff (BCH) or the Floquet-Magnus series blanes09 . We also point out and numerically demonstrate that in order to infer the correct behavior in the thermodynamic limit the Heisenberg time $t_{\rm H}$ (the inverse level spacing) has to be much larger than any relevant dynamical timescale in the system, specifically the long prethermalization time $T$. In particular, previous claims Heyl ; Polkovnikov ; Bukov of a small Trotter step transition (or a crossover) in system’s properties, like the scaling of errors, were a finite size effect caused by a finite averaging time (i.e., being in the incorrect limit $T>t_{\rm H}$). In the thermodynamic limit (TDL) there is no transition.
@@
Another utility of the truncated propagator is to study discrete time crystals (DTC) vedika16 ; else16 ; curt16 , see Ref. zaletel23 for a review. Because we have access to exact $T$ we will be able to address the question of the stability and the thermodynamic limit behavior of the DTC in a nearest-neighbor kicked Ising model away from its integrable $\tau=\pi$ point. We find that the prethermalization, underlying the robustness of DTC, is rather prevalent in the kicked Ising model and not limited only to small values of $\tau$. Finally, looking at the quasimomentum dependence of the spectral gap we will be able to calculate the diffusion constant, either during a finite but long prethermalization time, or in the Hamiltonian limit.
@@ II Observables and truncated operator propagator
Figure 2: Truncated propagator and Ruelle-Pollicott resonances. (A) Momentum-resolved operator propagator $M(k)$ is truncated to infinite system size operators with their density supported on at most $r$ sites. (B) Spectrum of non-unitary $M(k=0)$ lies inside a dashed unit circle. Shown are spectra at truncation $r=8$ and different gate duration $\tau$ in the kicked Ising model (11). Spectrum can be used to identify prethermalization (e.g. $\tau=0.8$) as well as a discrete time crystal ($\tau=3.0$). The largest eigenvalue (e.g., zoom-in inset with red arrow in the top middle frame) is shown in (C). (C) Dependence of the spectral gap $\Delta$ on $\tau$. Circles and crosses mark $\tau$ for which we show convergence with $r$, and correlation functions, respectively. (D) Autocorrelation function of energy $H_{0}$ (10) and magnetization $Z=\sum_{j}\sigma^{\rm z}_{j}$ (right). Dotted curve is exponential decay with the rate $\Delta$ from (C), showing that the inverse gap $T=1/\Delta$ gives the duration of the prethermalization plateau. In the thermodynamic limit $L\to\infty$ the plateau at large time disappears for any $\tau$ – there is no Trotterization transition with $\tau$.
To evaluate $C_{A}(t)$ we need the time evolved observable $A(t)=U_{\tau}^{-t}AU_{\tau}^{t}$, which can be written in the Hilbert space of operators as ${|A(t)\rangle\!\rangle}={\cal M}^{t}{|A\rangle\!\rangle}$, where ${\cal M}$ is a linear (super)operator propagator that propagates observables (operators) by one step (one $U_{\tau}$). With this notation the correlation function is simply
@@
where the Hilbert-Schmidt inner product is ${\langle\!\langle A|B\rangle\!\rangle}\equiv{\rm tr}(A^{\dagger}B)/{\rm tr}(\mathbbm{1})$. The expectation value is on the other hand $\langle A(t)\rangle={\langle\!\langle\mathbbm{1}|{\cal M}^{t}|A\rangle\!\rangle}$. In systems that have a mixing property (connected) correlation functions must decay to zero, while in chaotic ones without conserved operators one moreover expects this decay to be exponential for local $A$. One might wonder how can one obtain decaying correlation functions from the properties of ${\cal M}$? This is not obvious because ${\cal M}$ is unitary and therefore its spectrum is on a unit circle – one would expect that e.g. exponential decay of $C_{A}(t)$ will be according to Eq.(6) connected to eigenvalues of ${\cal M}$ that satisfy $|\lambda|<1$. Unitarity of ${\cal M}$ however is not in contradiction with the apparent relaxation (equilibration): namely, we expect relaxation of local observables and related decay of correlation functions; under unitary evolution no information is lost, it just gets spread out into the full Hilbert space, i.e., “disappears” in non-local operators.
In order to recover the decay of correlations from ${\cal M}$ one must look at analytical properties of its resolvent $({\cal M}-z\mathbbm{1})^{-1}$ inside a unit circle of complex $z$. This idea is known under the name of Ruelle-Pollicott (RP) resonances Ruelle86 ; Pollicott85 ; gaspard . One way to get access to possible singularities inside a unit circle is to limit oneself to an appropriate functional space – in our case it will be the space of local operators.
@@ II.1 The truncated propagator
The truncated operator propagator is the main technical tool of this work and will enable us to quantitatively study the shadowing, prethermalization, DTC, and more, thereby resolving several outstanding questions.
One way to break the unitarity of ${\cal M}$ is to truncate it to operators with local support. Doing that one obtains a finite dimensional truncated ${\cal M}$ whose eigenvalues $\lambda_{j}$ move inside a unit circle Prosen . The largest one in modulus $\lambda_{1}$ will, provided it is isolated from the rest, determine the asymptotic decay of correlations as $C_{A}(t)\sim\lambda_{1}^{t}$, and is called a Ruelle-Pollicott resonance Ruelle86 ; Pollicott85 , a resonance inside a unit circle in the analitical continuation of the propagator resolvent. In Ref. Prosen ${\cal M}$ was truncated to translationally invariant operators ($k=0$ in our notation) and studied in the kicked Ising model. Besides chaotic systems, the method can be also used to find constants of motion marcin , for instance in Ref. katja in cellular automata, and in Ref. U1 to discover that all magnetization conserving brickwall qubit circuits are integrable.
We shall use an extension of the above method to an arbitrary (quasi)momentum $k$ RP24 , a refinement that has been already used to find new unconventional symmetries in integrable circuits kpzwall . Momentum dependence will allow us to treat any observable, not just translationally invariant ones, as well as deduce transport properties, for instance the value of the diffusion constant. Namely, we are interested in translationally invariant $H_{0}$ without any spatial inhomogeneities, such that also its Trotterization $U_{\tau}$ will have homogeneous gates. In other words, $U_{\tau}$ commutes with a shift by one site operator ${\cal S}$ (the method can be generalized to systems invariant only under a shift by more than one site urban ), and therefore ${\cal M}$ is block diagonal in the eigenbasis of ${\cal S}$. Blocs can be labeled by eigenvalues ${\rm e}^{{\rm i}k}$ of ${\cal S}$, where $k\in[-\pi,\pi]$ is a lattice quasimomentum (we shall often call it just momentum). The operator basis in a given momentum block consists of operators $A^{(\bm{\alpha})}$,
@@
While it looks that in order to calculate $M(k)$ one has to deal with infinite-size operators $A^{(\bm{\alpha})}$, because of a sharp light-cone in quantum circuits the initial local density $a^{(\bm{\alpha})}_{j}$ can in one step spread over at most a fixed number $s$ of additional sites to the left or to the right of its initial support. For our specific circuit (Fig. 3) where $U_{\tau}$ contains only a single layer of commuting nearest-neighbor 2-qubit gates we have $s=1$. It is therefore enough to work with propagation on just $r+2$ sites. Denoting $b(\bm{\beta})\equiv U_{\tau}^{\dagger}a^{(\bm{\beta})}_{1}U_{\tau}$, Eq.(8) explicitly evaluates to
@@
Such an operator propagator has a symmetry $M^{\dagger}(U)=M(U^{\dagger})$ and is without truncation a unitary operator (for $k=0$ and Hermitian basis, like the Pauli basis (7), it is real orthogonal). If we truncate the basis to all local densities supported on at most $r$ consecutive sites, that is vector indices $\bm{\alpha}$ and $\bm{\beta}$ run over $3\times 4^{r-1}$ possible local densities ($3$ instead of $4$ comes because the first index $\alpha_{1}$ should not describe $\mathbbm{1}$, otherwise the density would be supported on at most $r-1$ sites), we get a truncated operator propagator that is the main workhorse of this work. Due to the truncation we are dropping some matrix elements of a unitary operator, resulting in a non-unitary matrix. For more details on implementation see Ref. RP24 ; urban .

### v2

@@ Abstract
One of premier utilities of present day noisy quantum computers is simulation of many-body quantum systems. We study how long in time is such a discrete-time simulation representative of a continuous time Hamiltonian evolution, namely, a finite time-step introduces so-called Trotterization errors. We show that the truncated operator propagator (Ruelle-Pollicott resonances) is a powerful tool to that end, as well as to study prethermalization and discrete time crystals, including finding those phenomena at large gate duration. We show that the effective energy is more stable than suggested by Trotter errors – a manifestation of prethermalization – while all other observables are not. Even the most stable observable though deteriorates in the thermodynamic limit. Different than in classical systems with the strongest chaos, where the faithfulness time (the shadowing time) can be infinite, in quantum many-body chaotic systems it is finite. A corollary of our results is also that, opposite to previous claims, there is no Trotterization transition in non-integrable many-body quantum systems. We demonstrate our results on a one-dimensional (1d) kicked Ising model, as well as on 1d kicked XX model and 2d kicked Ising model. The truncated propagator is also used to calculate the energy diffusion constant in the tilted-field Ising model with high accuracy.
@@ I Introduction
One of the defining property of chaotic systems Ott is exponential complexity of their time evolution as quantified by positive dynamical entropies. A manifestation of that is, for instance, the colloquial “butterfly effect” – two close initial conditions will separate exponentially fast in time. Considering that typical chaotic systems are not analytically solvable, and one therefore has to resort to the indignity of numerical simulations, a pertinent question is how representative are such noisy numerical trajectories of true trajectories? At first sight it would seem that the exponential separation of trajectories renders any noisy numerical simulation futile. However, things are not as grim.
First, we should specify what is it that we really want from a noisy simulation. Comparing a noisy and a true trajectory both starting from the same specific initial condition is not what we should in most cases care about. This is especially true in a many-body system, where we do not have access to all degrees of freedom, and in quantum mechanics where the stress is on (ensemble) averages rather than on individual pure states. For instance, calculating the overlap of two states, often called the fidelity (a.k.a. the Loschmidt echo), is not the most informative quantity in a many-body system (unless one is e.g. interested in the ground-state physics of a gapped system): small fidelity does not necessarily mean that few-body observables that we care about have gone awry. Furthermore, the properties of a specific exact initial state are really irrelevant; measurable quantities away from phase transitions should be robust to small variations. What one should care about is that the noisy evolution of a particular initial state is still representative of a true non-noisy evolution, i.e., that its results are similar to an exact average. It would be enough for the noisy evolution to be similar to a true evolution from some nearby initial condition. Precisely that property is captured by the famous shadowing property of chaotic trajectories Ott that comes to our rescue when doing noisy simulations of chaotic systems.
@@
It states that, while it is true that a trajectory $f^{t}_{\epsilon}(\mathbf{x}_{0})$ starting from $\mathbf{x}_{0}$ will in the presence of noise $\epsilon$ diverge exponentially fast from a true one $f^{t}_{0}(\mathbf{x}_{0})$ starting from $\mathbf{x}_{0}$, there exists a nearby initial condition $\mathbf{x}_{0}^{\prime}=\mathbf{x}_{0}+\bm{\delta}$ for which a true trajectory $f^{t}_{0}(\mathbf{x}_{0}^{\prime})$ closely follows the original noisy one, $f^{t}_{0}(\mathbf{x}_{0}^{\prime})\approx f^{t}_{\epsilon}(\mathbf{x}_{0})$. In other words, while a noisy trajectory is not a good representative of an exact one, it closely follows a “shadow trajectory” that starts from a slightly perturbed initial condition (Fig. 1). The shadowing time is the time up-to which one can find the shadow trajectory to follow the noisy one. For uniformly hyperbolic foothyper systems the above intuitive picture can be made rigorous in the form of the shadowing lemma Anosov67 ; Bowen75 , saying that the shadowing time in such systems is infinite. Uniformly hyperbolic systems are in a way the nicest chaotic systems (strongest chaos), an example being the Arnold’s cat map, in which noisy trajectories are representative of exact ensemble averages. For classical systems that are not uniformly hyperbolic the shadowing time is finite Yorke97 .
We address the question of the shadowing time in quantum many body systems. This is an important theoretical question, that is moreover very relevant for noisy quantum computers and simulators, whose best-case present use troyer23 is to simulate daley22 other quantum systems ibm ; google ; dario ; shtanko25 ; feig25 , including the prethermalization regime banuls23 that we shall discuss. If one wants to demonstrate the supremacy of quantum computation the number of qubits has to be large, and the simulation has to be faithful to true dynamics. Simulating many-body quantum Hamiltonian $H_{0}=\sum_{j}h_{j}$ with local interactions $h_{j}$ can be implemented by splitting the time evolution propagator
@@
Characterizing shadowing properties of a many-body quantum system is a formidable task, first because of all the intricacies that quantum mechanics brings, and second because of the system size limit $L\to\infty$, where even for classical systems things are not clear. We will therefore tackle a simpler question. We will not address the full shadowing, which is a statement on closeness of states, i.e., of all observables, but rather a question of direct experimental relevance that will nevertheless shed light also on the shadowing time. We will (i) focus on a specific type of noise, namely errors childs19 ; cubitt21 ; Heyl due to Trotterizing Hamiltonian evolution (2), and (ii) measure closeness by simple expectation values or autocorrelation functions of local observables (rather than some state norm).
I.1 Summary of results
@@
We will find that even in the optimal situation – looking at the best-behaving observable – a noisy simulation follows a true one only up-to a finite but long time, meaning that the shadowing time is necessarily finite. More specifically, under Trotterization errors there is a special observable, namely the energy $H_{0}$, that is very stable and resilient to errors, more than the simple Trotter error ${\cal O}(\tau^{2})$ would suggest. Despite its stability though, with increasing system size $L$ the stability will deteriorate. While in a finite system it might look that $H_{0}$ is stable, in the thermodynamic limit of $L\to\infty$ and time $t\to\infty$, no matter how small $\tau$, the energy $H_{0}$ will not be conserved implying that the noisy simulation is fundamentally different than the true one. As a consequence, perfect infinite-time shadowing, like the one holding for classical systems with the strongest chaos, is impossible under Trotterization errors in locally interacting quantum many-body systems. The relevant timescale of the above instability is nothing but the prethermalization Mori16 ; Abanin17 time $T$, for a review of prethermalization see Ref. ho23 , and is given by the largest eigenvalue of the momentum-resolved truncated operator propagator. As predicted theoretically, it is exponentially large in the gate frequency $1/\tau$, but the benefit of our approach will be to give an exact quantitative prediction for $T$, not just a bound. The corresponding eigenvector is on the other hand an almost conserved effective energy which is for small $\tau$ close to $H_{0}$. Note that the prethermalization time and the shadowing time are different concepts. The shadowing time is related to the worst-case stability – thermalization of observable that deviates first, whereas we are focusing on the best-case scenario, i.e., thermalization of $H_{0}$. An alternative viewpoint is that we are interested in the breaking of a single conserved operator $H_{0}$ due to Trotterization errors.
@@
The truncated propagator therefore offers a way to obtain the effective Hamiltonian useful in studies of Floquet systems Eckardt17 , and is an explicit and concrete alternative to various divergent operator expansions like the Baker-Cambell-Hausdorff (BCH) or the Floquet-Magnus series blanes09 . We also point out and numerically demonstrate that in order to infer the correct behavior in the thermodynamic limit the Heisenberg time $t_{\rm H}$ (the inverse level spacing) has to be much larger than any relevant dynamical timescale in the system, specifically the long prethermalization time $T$. It is another example when relying on small-system results, for instance obtained by exact diagonalization, can lead to incorrect conclusions. Previous claims Heyl ; Polkovnikov ; Bukov of a small Trotter step transition (or a crossover) in system’s properties, like the scaling of errors, were a finite size effect caused by a finite averaging time (i.e., being in the incorrect limit of $T>t_{\rm H}$). In the thermodynamic limit (TDL) there is no transition.
@@
Another utility of the truncated propagator is to study discrete time crystals (DTC) vedika16 ; else16 ; curt16 , see Ref. zaletel23 for a review. Because we have access to exact $T$ we will be able to address the question of the stability and the thermodynamic limit behavior of the DTC in a nearest-neighbor kicked Ising model away from its integrable $\tau=\pi$ point. We find that the prethermalization, underlying the robustness of DTC, is rather prevalent in the kicked Ising model and not limited only to small values of $\tau$. Finally, looking at the quasimomentum dependence of the spectral gap we will be able to calculate the diffusion constant, either during a finite but long prethermalization time, or in the Hamiltonian limit. Because we directly work in the thermodynamic limit $L\to\infty$ the value of the energy diffusion constant for the chaotic Ising model is calculated with higher precision that with most other methods. We also briefly show results for the kicked XX model and a two-dimensional (2D) kicked Ising model, demonstrating generality of our results.
@@ II Observables and truncated operator propagator
Figure 2: Truncated propagator and Ruelle-Pollicott resonances. (A) Momentum-resolved operator propagator $M(k)$ is truncated to infinite system size operators with their density supported on at most $r$ sites (sine-shaped curves suggest quasimomentum modulation of local densities, e.g., of $a^{(\bm{{\alpha}})}_{j}=\sigma^{\rm z}_{j}$ having $r=1$). (B) Spectrum of non-unitary $M(k=0)$ lies inside a dashed unit circle. Shown are spectra at truncation $r=8$ and different gate duration $\tau$ in the kicked Ising model (11). Spectrum can be used to identify prethermalization (e.g. $\tau=0.8$) as well as a discrete time crystal ($\tau=3.0$). The largest eigenvalue (e.g., zoom-in inset with red arrow in the top middle frame) is shown in (C). (C) Dependence of the spectral gap $\Delta$ on $\tau$. Circles and crosses mark $\tau$ for which we show convergence with $r$, and correlation functions, respectively. (D) Autocorrelation function of energy $H_{0}$ (10) and magnetization $Z=\sum_{j}\sigma^{\rm z}_{j}$ (right). Dotted curve is exponential decay with the rate $\Delta$ from (C), showing that the inverse gap $T=1/\Delta$ gives the duration of the prethermalization plateau. In the thermodynamic limit $L\to\infty$ the plateau at large time disappears for any $\tau$ – there is no Trotterization transition with $\tau$.
[…]
@@
[…]
@@ II.1 The truncated propagator
The truncated operator propagator is the main technical tool of this work and will enable us to quantitatively study the prethermalization, DTC, and more, thereby resolving several outstanding questions.
[…]
@@
While it looks that in order to calculate $M(k)$ one has to deal with infinite-size operators $A^{(\bm{\alpha})}$, because of a sharp light-cone in quantum circuits the initial local density $a^{(\bm{\alpha})}_{j}$ can in one step spread over at most a fixed number $s$ of additional sites to the left or to the right of its initial support. For our specific kicked Ising circuit (Fig. 3) where $U_{\tau}$ contains only a single layer of commuting nearest-neighbor 2-qubit gates we have $s=1$. It is therefore enough to work with propagation on just $r+2$ sites. One can in fact implement evolution of local operators in-situ using only $r$ sites instead of $r+2$, see Appendix A.2 for details, which enables us to study larger truncation sizes $r$, or simply speeding up calculations compared to tracking evolution on full $r+2$ sites. Denoting $b(\bm{\beta})\equiv U_{\tau}^{\dagger}a^{(\bm{{\beta}})}_{1}U_{\tau}$, Eq.(8) explicitly evaluates to
@@
Such an operator propagator has a symmetry $M^{\dagger}(U)=M(U^{\dagger})$ and is without truncation a unitary operator (for $k=0$ and Hermitian basis, like the Pauli basis (7), it is real orthogonal). If we truncate the basis to all local densities supported on at most $r$ consecutive sites (1d systems), that is vector indices $\bm{\alpha}$ and $\bm{\beta}$ run over $N=3\times 4^{r-1}$ possible local densities ($3$ instead of $4$ comes because the first index $\alpha_{1}$ should not describe $\mathbbm{1}$, otherwise the density would be supported on at most r1r-1 sites, and could be obtained by a translation of other basis elements), we get a truncated operator propagator that is the main workhorse of this work. Due to the truncation we are dropping some matrix elements of an infinite-size unitary operator, resulting in a finite non-unitary matrix of size $N\times N$. For more details on implementation see Ref. RP24 ; urban , and for in-situ propagation Appendix A.2.